Prove the trigonometric identity
2sinē(θ)-1 = sin⁴(θ)-cos⁴(θ)

Answer by Edwin McCravy(20060)

2sinē(θ)-1 = sin⁴(θ)-cos⁴(θ)
           = [sinē(θ)-cosē(θ)][sinē(θ)+cosē(θ)]
           = [sinē(θ)-cosē(θ)][1]
           = sinē(θ)-cosē(θ)
           = sinē(θ)-[1-sinē(θ)]
           = sinē(θ)-1+sinē(θ)
           = 2sinē(θ)-1 

Edwin